OpenStudy (anonymous):

Medal for best answer!!! −12 = 2 + 5v + 2v

OpenStudy (anonymous):

Simplify: -12 = 2 + 5v + 2v Add 5v + 2v: -12 = 2 + 7v Subtract 2 to both sides, what's -12 - 2?

OpenStudy (anonymous):

@Aleah.wood0613

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

Yep, that gives us -14 = 7v. Now divide 7 to both sides, what's -14 / 7?

OpenStudy (anonymous):

@Aleah.wood0613